OpenStudy (anonymous):

Algebra 2 practice exam?? Someone help me out??

OpenStudy (anonymous):

It's Lesson 1 Unit 8. Anyone from Connections?

hartnn (hartnn):

do you have any specific question ? if yes post it ..... or,are you asking for direct answers ? if yes, that is against our CoC here. http://openstudy.com/code-of-conduct please ask for specific questions so that we can help you learn, getting direct answers will not help you learn. and yes, giving direct ansers is also against CoC here.

OpenStudy (anonymous):

I never said give direct answers I was just wondering if anyone from Connections would help me out with a exam? Haha. I've already read the code of conduct before.

hartnn (hartnn):

so, its the 1st case, if you have, post specific questions... which topic exactly u need help with ?

OpenStudy (anonymous):

Math, and this is the question "Write the equation in slope-intercept form. What are the slope and y-intercept?" –3x – 8y = 10

hartnn (hartnn):

slope intercept form : y=mx+c [m is slope, and c is y-intercept] so, you need to isolate y from -3x-8y=10 can you ?

hartnn (hartnn):

add 8y to both sides, then subtract 10 from both sides..

OpenStudy (anonymous):

Wouldn't you subtract -3x instead of 10?

hartnn (hartnn):

subtract -3x ? no, we generally do inverse operation. since there is "-" (minus) with 3x, we ADD 3x to both sides, that will be another way to isolate y, you can try that also.

OpenStudy (anonymous):

I meant add I promise aha. I know how to do the inverse operations Soo We have -3x=18 Add -3x to both sides, and get x=21?

hartnn (hartnn):

where di y go ? :O -3x-8y=10 ADD 3x on both sides, -8y = 10+3x now, divide both sides by - 8....

hartnn (hartnn):

?